OpenStudy (anonymous):

2 power 56 when divided by 17 leaves a remainder of?

OpenStudy (anonymous):

\[2^{56}/17 = 4.239*10^{15}\] does that answer work for you or would you like more steps?

OpenStudy (anonymous):

This is same as 56*56 = 3136, when divided by 17 = 184 remainder 8

OpenStudy (anonymous):

2^16 congruent to 1 mod 17 so 2^56 = 2^8 mod17...

OpenStudy (anonymous):

56*56 nowhere near 2^56!!

OpenStudy (anonymous):

@DevinBlade remainder is a number less than 17

OpenStudy (anonymous):

oh so its 2^(56/17)? aha my bad

OpenStudy (anonymous):

\[\frac{2^{56}}{17}\]

OpenStudy (anonymous):

ohh the remainder x_x its been so long since i did old school division totally forgot about remainders

OpenStudy (anonymous):

I guess the answer is one, follow my next post to find why.

OpenStudy (anonymous):

@alsamixer Correct.

OpenStudy (anonymous):

\[2^{56} = 2^{8\times7} = (2^4)^{14} = (2^4 + 1 -1)^{14} = (a-1)^{14}\] where \[a = 2^{4}+1 = 17 \] Now expand the first equation using binomial equation. And all terms till the last term is divisible by 17, what remains is \[1^{14} = 1\] , hence the remainder is 1.

OpenStudy (anonymous):

If you don't know it you have a^(p-1) = 1 mod p (Fermat's Little Theorem).

OpenStudy (anonymous):

But does it apply to this problem? We wanted something like: \[(a-1)^p = 1 \mod a\]

OpenStudy (anonymous):

I posted it above already 2^16 congruent to 1 mod 17 so 2^56 = 2^8 mod17... 2^16 three times is 1 mod 17 so goes away and 2^8/17 is simple..